OpenStudy (anonymous):

Kendra has $5.25 in nickels and quarters. If she has 15 more quarters than nickels, how many of each coin does she have?

OpenStudy (anonymous):

let number of quarters be x so value in cents is 25x according to question number of nickels =x+15 so value of this is 5(x+15) now equation will be 25x + 5x+75 = 525 cents solve for x which is number of quarters plug in value of x in x+15 which will give you number of nickels @ElBrazilHue

OpenStudy (anonymous):

x=15 @aryandecoolest

OpenStudy (anonymous):

YEAH CORRECT SO NUMBER OF NICKELS WILL BE 15+15=30 :) @ElBrazilHue